The rate constant, k, for the first order reaction A --> 2B is 0.0424 s-1. If the initial concentration is 2.14M, what is the concentration after 56.114 seconds?
Answer #1

Answer-

Given,

Rate Constant = 0.0424 s-1

Initial Concentration = 2.14 M

Time = 56.114 s

Concentration at given time = ?

We know that,

where,

A = Concentration at time, t

A = Initial Concentration

k = rate constant

t = time

Put all the values,

[A] = 2.14 * e-(0.0424 * 56.114)

[A] = 2.14 * e-2.38

[A] = 2.14 * 0.0925

[A] = 0.19795 M

So, the concentration after 56.114 seconds is 0.198 M.